Find an equation of a parabola satisfying the given information. Focus (7,4), directrix x = - 8
Answer #1

öll Gimen focus (7,0) & Directix => x=-8 Tor parabola . (y-k)?= 4p(2-4) Focus= (htp, k) e Directory => x=h-p (hap, k) = (7,4)